OpenStudy (anonymous):

the length of a rectangle is three times its width. if the perimeter of a rectangle is 40m find its area

OpenStudy (anonymous):

2(3x + x) = 40 Solve for x then plug it in here: A = 3x^2

OpenStudy (anonymous):

so if 2 times 3 equals 5x what do i do then?

OpenStudy (anonymous):

oh wait x equals 10

OpenStudy (anonymous):

What? how did you get that? 2(3x + x) = 40 2(4x) = 40 8x = 40 x = 5 Can you finish up?

OpenStudy (anonymous):

i thought we had to distribute first but ok so 3x5x2= 30?

OpenStudy (anonymous):

It doesn't matter which you do first. 2(3x + x) = 40 6x + 2x = 40 8x = 40 x = 5 Then, plug it in: \[3x^2 \implies 3(5)^2 \implies ?\]

OpenStudy (anonymous):

75?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np :)